You own a portfolio that has 6,200 shares of stock A, which is priced at 15.5 dollars per share and has an expected return of 8.22 percent, and 1,000 shares of stock B, which is priced at 29.7 dollars per share and has an expected return of 16.15 percent. The risk-free return is 3.09 percent and inflation is expected to be 1.72 percent. What is the expected real return for your portfolio? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Weight of Stock A in Portfolio = (15.50*6,200/((15.50*6,200) + (29.70*1,000)))

Weight of Stock A in Portfolio = 0.7639

Weight of Stock B in Portfolio = 1 - 0.7639 = 0.2361

So,

Expected Return of Portfolio = 0.7639(0.0822) + 0.2361(0.1615)

Expected Return of Portfolio = 10.09%

Real Rate = (1 + Nominal Rate)/(1 + Inflation) - 1

Real Rate = (1.1009/1.0172) - 1

Real Rate = 8.23%